# Batched Nonparametric Contextual Bandits

### v3

@@ B.3.1 Controlling $V_{i}$
Fix some $11\leq i\leq M-$ 1, and some bin $C\in\mathcal{B}_{i}$. On the event $\mathcal{G}_{C}$ we have $\mathcal{I}_{\mathsf{p}(C)}^{{}^{\prime}}\subseteq\bar{\mathcal{I}}_{\mathsf{p}(C)}$, that is, for any $k\in\mathcal{I}_{\mathsf{p}(C)}^{{}^{\prime}}$,
@@
Here, step (i) uses relation (16), and the fact that $\pi_{t}(X_{t})\in\mathcal{I}_{\mathsf{p}(C)}^{{}^{\prime}}$ when $X_{t}\in C$. For step (ii), if $C$ is split, then it is no longer live, so the live regret incurred on the remaining batches is zero. On the other hand, if $C$ is not split, then $|\mathcal{I}_{C}^{\prime}|=1$. Without loss of generality, assume that arm $-1$ is eliminated. Conditioned on $\mathcal{A}_{C}$, this means $-1\notin\underline{\mathcal{I}}_{C}$ and there exists $x_{0}\in C$ such that $f^{(1)}(x_{0})-f^{(-1)}(x_{0})>c_{0}|C|^{\beta}$. By the smoothness condition, having a gap at least $c_{0}|C|^{\beta}$ on a single point in $C$ implies $f^{(1)}(x)-f^{(-1)}(x)>|C|^{\beta}$ for all $x\in C$. Therefore, arm 1 which is the remaining one is the optimal arm for all $x\in C$ and would not incur any regret further. The third inequality holds since $\mathbf{1}(\mathcal{G}_{C}\cap\mathcal{A}_{C})\leq 1$.
@@ B.3.2 Controlling $U_{i}$
Fix some $11\leq i\leq M-$ 1, and some bin $C\in\mathcal{B}_{i}$. Again, using the definition of $\mathcal{G}_{C}$, we obtain
@@ B.5.2 Proof of Lemma 8
Recall $U(\tau,T,C)=4\sqrt{\frac{\log(2T|C|^{d})}{\tau}}$ By relation (12), we can write
@@
    where step (i) uses the definition of ¯$\underline{\mathcal{I}}_{C}$. Consequently, ${|\bar{Y}_{C,i}^{(k^{\prime})}-\bar{f}_{C}^{(k^{\prime})}|\leq U(m_{C,i},T,C)/4}$ and $|\bar{Y}_{C,i}^{(k)}-\bar{f}_{C}^{(k)}|\leq U(m_{C,i},T,C)/4$ cannot hold simultaneously. Otherwise, this would contradict with $\bar{Y}_{C,i}^{(k^{\prime})}-\bar{Y}_{C,i}^{(k)}>U(m_{C,i},T,C)$ because $m_{C,i}\leq 2m_{C,i}^{\star}$ under $E$. Therefore,
@@
    Upper bounding $\mathbb{P}^{\mathcal{G}_{C}}((\mathcal{D}_{C}^{1})^{c}\cap\mathcal{D}_{C}^{2})$: when $(\mathcal{D}_{C}^{1})^{c}\cap\mathcal{D}_{C}^{2}$ happens, no arm in ¯$\underline{\mathcal{I}}_{C}$ is eliminated while some $k\notin\bar{\mathcal{I}}_{C}$ remains in the active arm set. By definition, there exists $x^{(k)}$ such that ${f^{\star}(x^{(k)})-f^{(k)}(x^{(k)})>8c_{0}|C|^{\beta}}.$ Let $\eta(k)$ be any arm that satisfies $f^{\star}(x^{(k)})=f^{(\eta(k))}(x^{(k)})$, and one can easily verify $\eta(k)\in\underline{\mathcal{I}}_{C}$. Since $k$ is not eliminated, we have $\bar{Y}_{C,i}^{(\eta(k))}-\bar{Y}_{C,i}^{(k)}\leq U(m_{C,i},T,C)$. On the other hand,
@@ Appendix C Proof of Theorem 4
To construct a hard instance, we partition $[0,1]$ into $z$ bins with equal width. Denote the bins by $C_{j}$ for $j=1,...,z$, and let $q_{j}$ be the center of $C_{j}$. Define a function $\phi:[0,1]\mapsto\mathbb{R}$ as $\phi(x)=(1-|x|)\mathbf{1}\{|x|\leq 1\}.$ Correspondingly define a function $\varphi_{j}:[0,1]\mapsto\mathbb{R}$ as $\varphi_{j}(x)=D_{\phi}z^{-1}\phi(2z(x-q_{j}))\mathbf{1}\{x\in C_{j}\},$ where $D_{\phi}=\min(2^{-1}L,1/4)$. Define a function $f:[0,1]\mapsto\mathbb{R}$:
@@ C.2 Three failure cases for $g$
$R_{T}(\hat{\pi})$ $\overset{\mathrm{(i)}}{\geq}\mathbb{E}\left[\sum_{t=t_{1}+1}^{t_{2}}\left(f^{\star}(X_{t})-f^{\hat{\pi}_{t}(X_{t})}(X_{t})\right)\right]$
=$\overset{\mathrm{(ii)}}{=}\mathbb{E}\left[\sum_{t=t_{1}+1}^{t_{2}}\left(f^{\star}(X_{t})-f^{\hat{\pi}_{t}(X_{t})}(X_{t})\right)\mathbf{1}\{X_{t}\in C_{1}\}\right]$
@@
$R_{T}(\hat{\pi})$ $\geq\sum_{t=t_{1}+1}^{t_{2}}\sum_{l=H/4}^{3H/4}\mathbb{E}\left[\left(f^{\star}(X_{t})-f^{\hat{\pi}_{t}(X_{t})}(X_{t})\right)\mathbf{1}\{X_{t}\in B_{l}\}\right]$
@@
$R_{T}(\hat{\pi})$ $\overset{\mathrm{}}{\geq}\mathbb{E}\left[\sum_{t=t_{1}+1}^{t_{2}}\left(f^{\star}(X_{t})-f^{\hat{\pi}_{t}(X_{t})}(X_{t})\right)\right]$
@@
$R_{T}(\hat{\pi})$ $\overset{\mathrm{}}{\geq}\mathbb{E}\left[\sum_{t=t_{2}+1}^{T}\left(f^{\star}(X_{t})-f^{\hat{\pi}_{t}(X_{t})}(X_{t})\right)\right]$
@@ Appendix D Proof of Theorem 5
Recall $\gamma(\alpha^{\star})=\beta(\alpha^{\star}+1)/(2\beta+d)=(\alpha^{\star}+1)/3=\gamma^{\star}$. Fix some small constant $\epsilon>0$. We first deal with $M=2$ and then extend the analysis to general $M$. Throughout the proof, we use the fact that the algorithm needs to provide its first batch size $t_{1}$ prior to the game, and design $\alpha^{\star}$ such that any choice of $t_{1}$ would fail.
@@ D.1 When $M=2$
Take $\alpha^{\star}=1$. Since fixed grid and adaptive grid are the same when $M=2$, by relation (9), we have
@@ D.2 General $M$
for $2\leq i\leq M$. Moreover, take $z_{2}=\lceil(36t_{1}(M-1)^{2})^{1/(2\beta+d)}\rceil$ and $z_{i}=\lceil(36T_{i-1}(M-1)^{2})^{1/(2\beta+d)}\rceil$ for $3\leq i\leq M$. Recall $Q_{i}(\cdot)=\sum_{\omega\in\Omega_{s_{i}}}q_{i}(\omega)\mathbb{P}_{\pi,i,\omega}(\cdot)$ from (13) given $z_{i}>0$ and $s_{i}\coloneqq\lceil z_{i}^{d-\alpha\beta}\rceil$. We can apply Lemma 3 with $k=2$ to obtain that there exists $i^{\star}\in\{2,...,M\}$ such that $Q_{i^{\star}}(A_{i^{\star}})\geq 1/(2(M-1))$. Consequently, by Lemma 4 we reach
$\sup_{(f,\frac{1}{2})\in\mathcal{F}(\alpha^{\star},\beta)}R_{T}(\pi;f)$ $\geq\sup_{f\in\mathcal{C}_{z_{i^{\star}}}}R_{T_{i^{\star}}}(\pi;f)$
$\apprge T_{i^{\star}}z_{i^{\star}}^{-\beta(1+\alpha^{\star})}(M-1)^{-\frac{1+\alpha^{\star}}{\alpha^{\star}}}$

### v4

@@ B.3.1 Controlling $V_{i}$
Fix some $1\leq i\leq M-1$, and some bin $C\in\mathcal{B}_{i}$. On the event $\mathcal{G}_{C}$ we have $\mathcal{I}_{\mathsf{p}(C)}^{{}^{\prime}}\subseteq\bar{\mathcal{I}}_{\mathsf{p}(C)}$, that is, for any $k\in\mathcal{I}_{\mathsf{p}(C)}^{{}^{\prime}}$,
@@
Here, step (i) uses relation (15), and the fact that $\pi_{t}(X_{t})\in\mathcal{I}_{\mathsf{p}(C)}^{{}^{\prime}}$ when $X_{t}\in C$. For step (ii), if $C$ is split, then it is no longer live, so the live regret incurred on the remaining batches is zero. On the other hand, if $C$ is not split, then $|\mathcal{I}_{C}^{\prime}|=1$. Without loss of generality, assume that arm 1-1 is eliminated. Conditioned on $\mathcal{A}_{C}$, this means $1\notin\underline{\mathcal{I}}_{C}$ and there exists $x_{0}\in C$ such that $f^{(1)}(x_{0})-f^{(-1)}(x_{0})>c_{0}|C|^{\beta}$. By the smoothness condition, having a gap at least $c_{0}|C|^{\beta}$ on a single point in $C$ implies $f^{(1)}(x)-f^{(-1)}(x)>|C|^{\beta}$ for all $x\in C$. Therefore, arm 1 which is the remaining one is the optimal arm for all $x\in C$ and would not incur any regret further. The third inequality holds since $\mathbf{1}(\mathcal{G}_{C}\cap\mathcal{A}_{C})\leq 1$.
@@ B.3.2 Controlling $U_{i}$
Fix some $1\leq i\leq M-1$, and some bin $C\in\mathcal{B}_{i}$. Again, using the definition of $\mathcal{G}_{C}$, we obtain
@@ B.5.2 Proof of Lemma 6
Recall $U(\tau,T,C)=4\sqrt{\frac{\log(2T|C|^{d})}{\tau}}$ . By relation (12), we can write
@@
    where step (i) uses the definition of $\underline{\mathcal{I}}_{C}$. Consequently, $|\bar{Y}_{C,i}^{(k^{\prime})}-\bar{f}_{C}^{(k^{\prime})}|\leq U(m_{C,i},T,C)/4$} and $|\bar{Y}_{C,i}^{(k)}-\bar{f}_{C}^{(k)}|\leq U(m_{C,i},T,C)/4$ cannot hold simultaneously. Otherwise, this would contradict with $\bar{Y}_{C,i}^{(k^{\prime})}-\bar{Y}_{C,i}^{(k)}>U(m_{C,i},T,C)$ because $m_{C,i}\leq 2m_{C,i}^{\star}$ under $E$. Therefore,
@@
    Upper bounding $\mathbb{P}^{\mathcal{G}_{C}}((\mathcal{D}_{C}^{1})^{c}\cap\mathcal{D}_{C}^{2})$: when $(\mathcal{D}_{C}^{1})^{c}\cap\mathcal{D}_{C}^{2}$ happens, no arm in $\underline{\mathcal{I}}_{C}$ is eliminated while some $k\notin\bar{\mathcal{I}}_{C}$ remains in the active arm set. By definition, there exists $x^{(k)}$ such that $f^{\star}(x^{(k)})-f^{(k)}(x^{(k)})>8c_{0}|C|^{\beta}$}. Let $\eta(k)$ be any arm that satisfies $f^{\star}(x^{(k)})=f^{(\eta(k))}(x^{(k)})$, and one can easily verify $\eta(k)\in\underline{\mathcal{I}}_{C}$. Since $k$ is not eliminated, we have $\bar{Y}_{C,i}^{(\eta(k))}-\bar{Y}_{C,i}^{(k)}\leq U(m_{C,i},T,C)$. On the other hand,
@@ Appendix C Proof of Theorem 4
To construct a hard instance, we partition $[0,1]$ into $z$ bins with equal width. Denote the bins by $C_{j}$ for j=1,,zj=1,...,z, and let $q_{j}$ be the center of $C_{j}$. Define a function $\phi:[0,1]\mapsto\mathbb{R}$ as $\phi(x)=(1-|x|)\mathbf{1}\{|x|\leq 1\}.$ Correspondingly define a function $\varphi_{j}:[0,1]\mapsto\mathbb{R}$ as $\varphi_{j}(x)=D_{\phi}z^{-1}\phi(2z(x-q_{j}))\mathbf{1}\{x\in C_{j}\},$ where $D_{\phi}=\min(2^{-1}L,1/4)$. Define a function $f:[0,1]\mapsto\mathbb{R}$:
@@ C.2 Three failure cases for $g$
[…]
$\overset{\mathrm{(ii)}}{=}\mathbb{E}\left[\sum_{t=t_{1}+1}^{t_{2}}\left(f^{\star}(X_{t})-f^{\hat{\pi}_{t}(X_{t})}(X_{t})\right)\mathbf{1}\{X_{t}\in C_{1}\}\right]$
@@
[…]
@@
[…]
@@
[…]
@@ Appendix D Proof of Theorem 5
Recall $\gamma(\alpha^{\star})=\beta(\alpha^{\star}+1)/(2\beta+d)=(\alpha^{\star}+1)/3=\gamma^{\star}$. Denote by $h_{M}(\gamma^{\star})=(1-\gamma^{\star})/(1-(\gamma^{\star})^{M})$. Fix some small constant $\epsilon>0$. We first deal with $M=2$ and then extend the analysis to general $M$. Throughout the proof, we use the fact that the algorithm needs to provide its first batch size $t_{1}$ prior to the game, and design $\alpha^{\star}$ such that any choice of $t_{1}$ would fail.
@@ D.1 When $M=2$
Take $\alpha^{\star}=1$. By relation (9), we have
@@ D.2 General $M$
for $2\leq i\leq M$. Moreover, take $z_{2}=\lceil(36t_{1}(M-1)^{2})^{1/(2\beta+d)}\rceil$ and $z_{i}=\lceil(36T_{i-1}(M-1)^{2})^{1/(2\beta+d)}\rceil$ for $3\leq i\leq M$. By relation (A.2) , one has
$\sup_{(f,\frac{1}{2})\in\mathcal{F}(\alpha^{\star},\beta)}R_{T}(\pi;f)$ $\apprge M^{-\frac{(\alpha\beta+1)(1+\alpha)}{\alpha}}T_{i}z_{i}^{-\beta(1+\alpha)}\left(\mathbb{E}_{\omega\sim\mathrm{Unif}(\Omega)}\mathbb{P}_{\pi,\omega}(A_{m})-\frac{1}{2(M-1)}\right)^{\frac{1+\alpha}{\alpha}},$
for any $2\leq i\leq M$. Since there exists $i^{\star}\in\{2,...,M\}$ such that $\mathbb{E}_{\omega\sim\mathrm{Unif}(\Omega)}\mathbb{P}_{\pi,\omega}(A_{i^{\star}})\geq 1/(M-1)$, we reach
$\sup_{(f,\frac{1}{2})\in\mathcal{F}(\alpha^{\star},\beta)}R_{T}(\pi;f)$ $\apprge M^{-\frac{(\alpha\beta+1)(1+\alpha)}{\alpha}}T_{i^{\star}}z_{i^{\star}}^{-\beta(1+\alpha)}\left(\mathbb{E}_{\omega\sim\mathrm{Unif}(\Omega)}\mathbb{P}_{\pi,\omega}(A_{i^{\star}})-\frac{1}{2(M-1)}\right)^{\frac{1+\alpha}{\alpha}}$
$\apprge M^{-\frac{(\alpha\beta+2)(1+\alpha)}{\alpha}}T_{i^{\star}}z_{i^{\star}}^{-\beta(1+\alpha)}$